# Contrastive Learning with Synthetic Positives

### v1

@@ 1 Introduction
This paper proposes a new approach called Contrastive Learning with Synthetic Positives (CLSP), which integrates additional synthetic positives into conventional contrastive training schema. These synthetic positives exhibit rich deformations and diverse backgrounds, allowing the model to emphasize generic features beyond naive data augmentation. Specifically, our method employs the unconditional diffusion model to generate positives. It has been proved that well-trained unconditional diffusion models can capture semantic information at their intermediate layers [4, 49]. During a random reversed diffusion process, we can replace the features of the intermediate layers with the semantic features extracted from an anchor image. This results in the generation of images possessing similar semantic content to the anchor image but differing in background and context due to the randomness of features in other layers. We consider these synthetic samples as “hard” positives of the anchor image because the semantic similarity could be subtle and extremely distorted. To learn from such positive pairs, we utilize an additional loss term, similar to [20], to minimize their feature similarity. Given the inherent slowness of the diffusion sampling process, generating synthetic positives on the fly is impractical, even with speedup solutions like DDIM [45]. Consequently, we pre-generate a positive candidate set with a size of $k$ ($k\leq 8$) for each sample in the training set. During training, we randomly sample one image from this set as the additional positive. Experimental results show that CLSP can significantly improve the representation learning performance when compared to its conventional counterpart.
@@ 2 Related Work
Contrastive Learning. Contrastive learning, an effective SSL approach, has shown remarkable success in various domains such as computer vision [34, 38, 56, 40], natural language processing [1, 52, 47] and graph learning [54, 21], rivaling supervised learning models. Various frameworks have been proposed to learn visual representations using contrastive learning. CPC [37] predicts the future output of sequential data by using a probabilistic contrastive loss to capture maximally useful information. SimCLR [7, 8] uses other samples from the mini-batch as negative pairs. MoCo [22, 9] uses a momentum-updated memory bank to maintain numerous negative samples. Given the necessity of learning from both positive and negative pairs, researchers have proposed several techniques to construct more informative positive and negative pairs to increase training efficiency. For example, [26, 32, 28, 44] apply Mixup [57] to generate hard negatives. [42] proposed a negative sampling strategy to reweight the contrastive learning objective. UnReMix [46] combines anchor similarity, model uncertainty, and representativeness to sample hard negative, which improves the results on several visual, text, and graph benchmark datasets. Other works use nearest neighbor to select multi-instance positives to increase the feature diversity within the same semantic class [17, 29, 18]. However, the nearest neighbor algorithm can only identify easy positives, which makes the improvement marginal when the discrimination task becomes harder. [48] address this challenge by jointly learning the contrastive learning model and a GAN, a pair of momentum-updated generators are proposed to generate hard positive pairs. However, the joint learning schema is highly unstable and hard to optimize. Other enhancements in contrastive learning involve optimizing the contrastive loss [11, 50, 53] or eliminating negative pairs such as SwaV [6], SimSiam [10], BYOL [20], and BarlowTwins [55]. These methods don’t face false negative issues, as they solely require positive pairs. Our work in this paper is built upon standard contrastive learning (e.g., SimCLR and MoCo) and focuses on the hard positive generation.
@@ 6 Discussion
Alternative Positive Generation Methods. While we demonstrated that using feature interpolation in diffusion sampling is an effective way to generate harder positives for contrastive learning, alternative methods exist for positive generation with diffusion models. For example, recent work RCG [33] proposes an approach where the diffusion model is conditioned on the embedding generated from the anchor image to generate high-quality images similar to the anchor image. However, this method requires a pre-trained encoder, usually trained with SSL, to generate embeddings as the condition. It remains unclear whether the generated positive could, in return, benefit the original SSL performance. In addition, we could cluster the embeddings into a fixed number of clusters and use the cluster as the condition. The sampling will be similar to the class-conditioned diffusion model, and the generated images will be more diverse because they are not directly related to the original embedding. We leave the further discussion and comparison in the supplementary.
Liminations. Despite the promising results achieved by our CLSP method, we identify the following limitations. (1) Computation and Memory Overhead. The additional computation costs of CLSP include the diffusion model training and sampling, which could be huge if the training set is large. Because we pre-generate the additional positives into a candidate set, additional memory is needed to save the generated positives. Besides, one additional branch is added to SimCLR, so an extra 50% computation overhead is needed. (2) Increased Number of Hyperparameters The positive generation part has two hyperparameters including the feature interpolation weight and candidate set size. The CLSP loss introduces additional loss weight as an extra hyperparameter.
@@ 0.A.1 Alternative Positive Generation Methods
In this section, we discuss two alternative positive generation methods using the diffusion model and compare them with our proposed CLSP. (1) Similar to the idea in [33], we use an image encoder (pre-trained by any SSL method) to transfer the raw image distribution to a low-dimensional semantic embedding. Subsequently, we train a conditional diffusion model to map a noise distribution to the image distribution conditioned on the semantic embedding. This approach, termed RCG, is illustrated in Fig. 0.A.1(a). Due to the influence of semantic embedding on the diffusion sampling process, the resulting synthetic images typically contain similar semantic content to that of the anchor image and can be used as additional positives. However, one of the key drawbacks of RCG is that the synthetic positives often closely resemble the anchor image, potentially diminishing the benefits of learning these "easy" positives compared to CLSP. As can be seen in Fig. 0.A.2, RCG-generated images exhibit a higher visual similarity to the anchor images, with less variation in semantic content. (2) To mitigate the aforementioned limitation in RCG, we can use a cluster head (e.g., MLP) to cluster the semantic embeddings generated in RCG to distinct clusters and then use these clusters as the conditional to guide diffusion sampling. We name this method RCG-cluster (Fig. 0.A.1(b)). The idea is inspired by the class conditional diffusion model [24], where a pre-trained model can generate diverse images belonging to a specific class given its label. In RCG-cluster, the cluster head is trained using unsupervised learning techniques such as k-means, and the resulting clusters serve as conditions similar to those in the class conditional diffusion model. RCG-cluster has the potential to generate even more diverse positives than CLSP because the generated images are not explicitly correlated with the semantic embedding of the anchor image. However, it may introduce an increased risk of false positives when the anchor image is not mapped into the right cluster (Fig. 0.A.2).
@@ 0.A.3 Visualization of Representations
We compare the representations learned by our proposed CLSP-SimCLR and standard SimCLR using t-SNE visualization in Fig. 0.A.3. It can be seen that SimCLR has more overlapped representations among different classes. Such overlapped representations are non-discriminative and offer less information for the downstream tasks. However, the clusters learned by the proposed CLSP-SimCLR are more dense and separable, meaning that the SSL model learns better discriminative features of different classes and provides more information for downstream tasks.

### v2

@@ 1 Introduction
This paper proposes a new approach called Contrastive Learning with Synthetic Positives (CLSP), which integrates additional synthetic positives into a conventional contrastive training schema. These synthetic positives exhibit rich deformations and diverse backgrounds, allowing the model to emphasize generic features beyond naive data augmentation. Specifically, our method employs the unconditional diffusion model to generate positives. It has been proved that well-trained unconditional diffusion models can capture semantic information at their intermediate layers [4, 49]. During a random reversed diffusion process, we can replace the features of the intermediate layers with the semantic features extracted from an anchor image. This results in the generation of images possessing similar semantic content to the anchor image but differing in background and context due to the randomness of features in other layers. We consider these synthetic samples as “hard” positives of the anchor image because the semantic similarity could be subtle and extremely distorted. To learn from such positive pairs, we utilize an additional loss term, similar to [20], to minimize their feature similarity. Given the inherent slowness of the diffusion sampling process, generating synthetic positives on the fly is impractical, even with speedup solutions like DDIM [45]. Consequently, we pre-generate a positive candidate set with a size of $k$ ($k\leq 8$) for each sample in the training set. During training, we randomly sample one image from this set as an additional positive. Experimental results show that CLSP can significantly improve the representation learning performance when compared to its conventional counterpart.
@@ 2 Related Work
Contrastive Learning. Contrastive learning, an effective SSL approach, has shown remarkable success in various domains such as computer vision [34, 38, 56, 40], natural language processing [1, 52, 47], and graph learning [54, 21], rivaling supervised learning models. Various frameworks have been proposed to learn visual representations using contrastive learning. CPC [37] predicts the future output of sequential data by using a probabilistic contrastive loss to capture maximally useful information. SimCLR [7, 8] uses other samples from the mini-batch as negative pairs. MoCo [22, 9] uses a momentum-updated memory bank to maintain numerous negative samples. Given the necessity of learning from both positive and negative pairs, researchers have proposed several techniques to construct more informative positive and negative pairs to increase training efficiency. For example, [26, 32, 28, 44] apply Mixup [57] to generate hard negatives. [42] proposed a negative sampling strategy to reweight the contrastive learning objective. UnReMix [46] combines anchor similarity, model uncertainty, and representativeness to sample hard negatives, which improves the results on several visual, text, and graph benchmark datasets. Other works use nearest neighbor to select multi-instance positives to increase the feature diversity within the same semantic class [17, 29, 18]. However, the nearest neighbor algorithm can only identify easy positives, which makes the improvement marginal when the discrimination task becomes harder. [48] address this challenge by jointly learning the contrastive learning model and a GAN, a pair of momentum-updated generators are proposed to generate hard positive pairs. However, the joint learning schema is highly unstable and hard to optimize. Other enhancements in contrastive learning involve optimizing the contrastive loss [11, 50, 53] or eliminating negative pairs such as SwaV [6], SimSiam [10], BYOL [20], and BarlowTwins [55]. These methods don’t face false negative issues, as they solely require positive pairs. Our work in this paper is built upon standard contrastive learning (e.g., SimCLR and MoCo) and focuses on the hard positive generation.
@@ 6 Discussion
Alternative Positive Generation Methods. While we demonstrated that using feature interpolation in diffusion sampling is an effective way to generate harder positives for contrastive learning, alternative methods exist for positive generation with diffusion models. For example, recent work RCG [33] proposes an approach where the diffusion model is conditioned on the embedding generated from the anchor image to generate high-quality images similar to the anchor image. However, this method requires a pre-trained encoder, usually trained with SSL, to generate embeddings as the condition. It remains unclear whether the generated positive could, in return, benefit the original SSL performance. In addition, we could cluster the embeddings into a fixed number of clusters and use the cluster as the condition. The sampling will be similar to the class-conditioned diffusion model, and the generated images will be more diverse because they are not directly related to the original embedding. We leave the further discussion and comparison to the supplementary.
Limitations. Despite the promising results achieved by our CLSP method, we identify the following limitations. (1) Computation and Memory Overhead. The additional computation costs of CLSP include the diffusion model training and sampling, which could be huge if the training set is large. Because we pre-generate the additional positives into a candidate set, additional memory is needed to save the generated positives. Besides, one additional branch is added to SimCLR, so an extra 50% computation overhead is needed. (2) Increased Number of Hyperparameters The positive generation part has two hyperparameters, including the feature interpolation weight and candidate set size. The CLSP loss introduces additional loss weight as an extra hyperparameter.
@@ 0.A.1 Alternative Positive Generation Methods
In this section, we discuss two alternative positive generation methods using the diffusion model and compare them with our proposed CLSP. (1) Similar to the idea in [33], we use an image encoder (pre-trained by any SSL method) to transfer the raw image distribution to a low-dimensional semantic embedding. Subsequently, we train a conditional diffusion model to map a noise distribution to the image distribution conditioned on the semantic embedding. This approach, termed RCG, is illustrated in Fig. 0.A.1(a). Due to the influence of semantic embedding on the diffusion sampling process, the resulting synthetic images typically contain similar semantic content to that of the anchor image and can be used as additional positives. However, one of the key drawbacks of RCG is that the synthetic positives often closely resemble the anchor image, potentially diminishing the benefits of learning these "easy" positives compared to CLSP. As can be seen in Fig. 0.A.2, RCG-generated images exhibit a higher visual similarity to the anchor images, with less variation in semantic content. (2) To mitigate the aforementioned limitation in RCG, we can use a cluster head (e.g., MLP) to cluster the semantic embeddings generated in RCG into distinct clusters and then use these clusters as the condition to guide diffusion sampling. We name this method RCG-cluster (Fig. 0.A.1(b)). The idea is inspired by the class conditional diffusion model [24], where a pre-trained model can generate diverse images belonging to a specific class given its label. In RCG-cluster, the cluster head is trained using unsupervised learning techniques such as k-means, and the resulting clusters serve as conditions similar to those in the class conditional diffusion model. RCG-cluster has the potential to generate even more diverse positives than CLSP because the generated images are not explicitly correlated with the semantic embedding of the anchor image. However, it may introduce an increased risk of false positives when the anchor image is not mapped into the right cluster (Fig. 0.A.2).
@@ 0.A.3 Visualization of Representations
We compare the representations learned by our proposed CLSP-SimCLR and standard SimCLR using t-SNE visualization in Fig. 0.A.3. It can be seen that SimCLR has more overlapped representations among different classes. Such overlapped representations are non-discriminative and offer less information for the downstream tasks. However, the clusters learned by the proposed CLSP-SimCLR are denser and separable, meaning that the SSL model learns better discriminative features of different classes and provides more information for downstream tasks.